# The Gray Zone of Faithfulness: Taming Ambiguity in Unfaithfulness Detection

### v3

@@ Abstract
Ensuring that Large Language Models (LLMs) generate summaries faithful to a given source document is essential for real-world applications. While prior research has explored LLM faithfulness, existing benchmarks suffer from annotation ambiguity, primarily due to the ill-defined boundary of permissible external knowledge in generated outputs. For instance, common sense is often incorporated into responses and labeled as “faithful”, yet the acceptable extent of such knowledge remains unspecified, leading to inconsistent annotations. To address this issue, we propose a novel faithfulness annotation framework, which introduces an intermediate category, Out-Dependent, to classify cases where external knowledge is required for verification. Using this framework, we construct VeriGray [^1] https://huggingface.co/datasets/Ding-Qiang/veri-gray (Verification with the Gray Zone) – a new unfaithfulness detection benchmark in summarization. Statistics reveal that even SOTA LLMs, such as GPT-5, exhibit hallucinations ($\sim 6\%$ of sentences) in summarization tasks. Moreover, a substantial proportion ($\sim 8\%$ on average of models) of generated sentences fall into the Out-Dependent category, underscoring the importance of resolving annotation ambiguity in unfaithfulness detection benchmarks. Experiments demonstrate that our benchmark poses significant challenges to multiple baseline methods, indicating considerable room for future improvement.
@@ 4.2 Human Annotation
Each sentence in the generated summaries was evaluated for faithfulness by two human annotators, following the annotation framework outlined in Figure 1. Specifically, if a sentence was not directly supported by the source document, annotators further assessed whether it could be verified using a combination of the document and external knowledge. External knowledge was retrieved via Bing Search. In cases where a sentence was labeled as Out-Dependent, annotators were required to cite URLs from reliable external sources—such as Wikipedia or mainstream media websites. Conversely, if a sentence was labeled as Unfaithful, annotators had to provide evidence from the source document along with a clear reasoning process. To ensure reproducibility, all relevant web pages referenced during the annotation process were saved and included as attachments [^2] They are also open sourced at https://huggingface.co/datasets/Ding-Qiang/veri-gray. to the dataset.
@@ 4.3 Dataset Analysis
Here, we investigate the distribution of unfaithfulness classes in our benchmark [^4] We also explore the sequential dependency of unfaithfulness in Appendix C to see if the phenomenon of hallucination snwoballing (zhang2024language) exists in our benchmark.. Figure 2 shows the sentence-level category breakdown across different models. Overall, the majority are Faithful (76%), with Explicitly-Supported sentences (63%) substantially outnumbering Implicitly-Supported sentences (13%). Unfaithful sentences constitute approximately 14% of the total, most of which are Fabricated (11%). A notable proportion of sentences fall under the Not Sure category, primarily as Out-Dependent (8%), while fewer than 1% are labeled as Ambiguous. Across models, DeepSeek-V3-0324 and GPT-5 yield the highest proportion of Not Sure sentences (15%). Meanwhile, Claude-3.5-Sonnet generates the fewest Unfaithful sentences (3%), whereas Qwen3-8B generates the most (29%), reflecting the comparative success of proprietary models in maintaining faithfulness.
@@ 5.1 Baselines
Zero-shot LLMs. Zero-shot LLMs are prompted with the definitions from our taxonomy (see Appendix E for the prompt). The evaluated LLMs include GPT-5, DeepSeek-R1 (guo2025deepseekr1), DeepSeek-V3, Qwen3 235B A22B (yang2025qwen3), and QwQ-32B (team2025qwq). We also included GPT-5 enhanced with RAG (lewis2020retrieval) – hereafter referred to as GPT-5 + RAG – to provide the LLM with access to external world knowledge and help verify Out-Dependent examples. The retrieval corpus was the webpages collected during annotation (see Section 4.2). For more details, please refer to Appendix E.
Others. The other baselines include two model-internals-based methods, LLM-Check (sriramanan2024llmcheck) and ReDeEP (sun2024redeep), a probability-based method CCP (Claim Conditioned Probability) (fadeeva2024factchecing), and a retrieve-then-verify approach InFusE (zhang2024finegrained). Implementation details for all baselines are provided in Appendix D.
@@ 5.2 Evaluation Protocol
To fully utilize the fine-grained annotations, we propose another protocol that evaluates the ranking quality of the hallucination detectors. Without filtering or merging labels, we consider the full dataset be $\{(x_{i},y_{i})\}_{i=1}^{n}$, where $y_{i}$ is the fine-grained label of instance $x_{i}$. For predictions, instead, the output space is merged into Faithful/Unfaithful as before, assigned with a faithfulness degree order of Unfaithful $\prec$ Faithful. Let $\mathcal{P}:=\{(x_{i},x_{j})|y_{i}\succ y_{j}\}$. Inspired by the ranking expression of AUC (Area Under the Curve; Eq. (2.21) in zhou2021machine), we propose a novel metric ranking loss:
@@ 5.3 Results
DeepSeek-V3 61.2 23.0 86.8 36.4 43.4
Qwen3 235B A22B 68.9 40.1 76.7 52.6 39.7
QwQ 32B 72.4 47.4 76.8 58.6 38.0
DeepSeek-R1 78.9 64.8 62.8 63.8 33.7
GPT-5 80.0 67.9 61.1 64.4 33.0
GPT-5 + RAG 81.4 69.3 66.1 67.7 32.6
@@
ANAH-v2 53.0 11.5 27.7 16.3 46.6
MiniCheck 7B 61.8 27.9 55.2 37.0 42.8
Lynx 8B Instruct v1.1 62.6 38.7 34.6 36.5 42.3
Lynx 70B Instruct 67.2 40.8 54.4 46.6 42.0
@@
LLM-Check (Attn Score) 52.2 56.8 16.7 25.8 48.0
LLM-Check (Hidden Score) 54.9 56.1 18.3 27.6 48.7
ReDeEP 57.6 67.9 19.2 30.0 43.0
InFusE 46.6 37.3 13.5 19.8 51.9
CCP 58.6 47.7 22.4 30.5 44.1
@@
The main evaluation results are shown in Table 5.3. As shown, the most effective methods are zero-shot LLMs, notably GPT-5 + RAG, which achieve balanced accuracy of over 80%, hallucination F1 of nearly 68%, and ranking loss of less than 33%. Figure 4 shows the class-wise precision and recall of zero-shot LLMs. Across all models, recalls for Ambiguous and Out-Dependent remain consistently low. Nevertheless, compared to GPT-5, GPT-5 + RAG significantly boosts the precision of Out-Dependent while maintaining the recall, showing the effectiveness of using RAG. The low recall for Ambiguous may stem from both the scarcity of Ambiguous examples and the models’ lack of sensitivity to linguistic ambiguity. As for Out-Dependent, we found this class was mainly misclassified as faithful classes (see Figure 7 in Appendix F), indicating that the low recall may be due to the LLM detectors’ unawareness of using external knowledge from their parametric memory.
@@ Appendix B Details of Generating the Summaries
Appendix C Sequential Dependency of Unfaithfulness
@@
Appendix D Implementation Details of Baselines
@@
Appendix E Zero-shot Detector Details
@@
Appendix F The Confusion Matrix of GPT-5 + RAG

### v4

@@ Abstract
Ensuring that Large Language Models (LLMs) generate summaries faithful to a given source document is essential for real-world applications. While prior research has explored LLM faithfulness, existing benchmarks suffer from annotation ambiguity, primarily due to the ill-defined boundary of permissible external knowledge in generated outputs. For instance, common sense is often incorporated into responses and labeled as “faithful”, yet the acceptable extent of such knowledge remains unspecified, leading to inconsistent annotations. To address this issue, we propose a novel faithfulness annotation framework, which introduces an intermediate category, Out-Dependent, to classify cases where external knowledge is required for verification. Using this framework, we construct VeriGray [^1] https://huggingface.co/datasets/Ding-Qiang/veri-gray (Verification with the Gray Zone) – a new unfaithfulness detection benchmark in summarization. Statistics reveal that even SOTA LLMs, such as GPT-5, exhibit hallucinations ($\sim 6\%$ of sentences) in summarization tasks. Moreover, a substantial proportion ($\sim 9\%$ on average of models) of generated sentences fall into the Out-Dependent category, underscoring the importance of resolving annotation ambiguity in unfaithfulness detection benchmarks. Experiments demonstrate that our benchmark poses significant challenges to multiple baseline methods, indicating considerable room for future improvement.
@@ 4.2 Human Annotation
Each sentence in the generated summaries was evaluated for faithfulness by two human annotators, following the annotation framework outlined in Figure 1. Specifically, if a sentence was not directly supported by the source document, annotators further assessed whether it could be verified using a combination of the document and external knowledge. External knowledge was retrieved via Bing Search. In cases where a sentence was labeled as Out-Dependent, annotators were required to cite URLs from reliable external sources—such as Wikipedia or mainstream media websites. Conversely, if a sentence was labeled as Unfaithful, annotators had to provide evidence from the source document along with a clear reasoning process. To ensure reproducibility, all relevant web pages referenced during the annotation process were saved and included as attachments [^2] They are also open sourced at https://huggingface.co/datasets/Ding-Qiang/veri-gray. to the dataset. For more details of the annotation guidelines, please refer to Appendix C.
@@ 4.3 Dataset Analysis
Here, we investigate the distribution of unfaithfulness classes in our benchmark [^4] We also explore the sequential dependency of unfaithfulness in Appendix D to see if the phenomenon of hallucination snwoballing (zhang2024language) exists in our benchmark.. Figure 2 shows the sentence-level category breakdown across different models. Overall, the majority are Faithful (74%), with Explicitly-Supported sentences (61%) substantially outnumbering Implicitly-Supported sentences (13%). Unfaithful sentences constitute approximately 15% of the total, most of which are Fabricated (12%). A notable proportion of sentences fall under the Not Sure category, primarily as Out-Dependent (9%), while fewer than 1% are labeled as Ambiguous. Across models, DeepSeek-V3-0324 and GPT-5 yield the highest proportion of Not Sure sentences (16%). Meanwhile, Claude-3.5-Sonnet generates the fewest Unfaithful sentences (3%), surpassing the more recent GPT-5 (6%), whereas Qwen3-8B generates the most (32%), reflecting the comparative success of proprietary models in maintaining faithfulness.
@@ 5.1 Baselines
Zero-shot LLMs. Zero-shot LLMs are prompted with the definitions from our taxonomy (see Appendix F for the prompt). The evaluated LLMs include GPT-5, DeepSeek-R1 (guo2025deepseekr1), DeepSeek-V3, Qwen3 235B A22B (yang2025qwen3), and QwQ-32B (team2025qwq). We also included GPT-5 enhanced with RAG (lewis2020retrieval) – hereafter referred to as GPT-5 + RAG – to provide the LLM with access to external world knowledge and help verify Out-Dependent examples. The retrieval corpus was the webpages collected during annotation (see Section 4.2). For more details, please refer to Appendix F.
Others. The other baselines include two model-internals-based methods, LLM-Check (sriramanan2024llmcheck) and ReDeEP (sun2024redeep), a probability-based method CCP (Claim Conditioned Probability) (fadeeva2024factchecing), and a retrieve-then-verify approach InFusE (zhang2024finegrained). Implementation details for all baselines are provided in Appendix E.
@@ 5.2 Evaluation Protocol
To fully utilize the fine-grained annotations, we propose another protocol that evaluates the ranking quality of the hallucination detectors. We consider the full dataset filtering out No-Fact instances only, denoted as $\mathcal{D}_{\text{fact}}=\{(x_{i},y_{i})|y_{i}\neq\texttt{No-Fact}\}$, where $y_{i}$ is the fine-grained label of instance $x_{i}$. For predictions, instead, the output space is merged into Faithful/Unfaithful as before, assigned with a faithfulness degree order of Unfaithful $\prec$ Faithful. Let $\mathcal{P}:=\{(x_{i},x_{j})|y_{i}\succ y_{j},(x_{i},y_{i}),(x_{j},y_{j})\in\mathcal{D}_{\text{fact}}\}$. Inspired by the ranking expression of AUC (Area Under the Curve; Eq. (2.21) in zhou2021machine), we propose a novel metric ranking loss:
@@ 5.3 Results
DeepSeek-V3 60.4 21.3 90.1 34.5 44.3
Qwen3 235B A22B 68.5 39.0 80.1 52.5 40.0
QwQ 32B 73.1 48.0 84.7 61.3 37.5
DeepSeek-R1 79.3 64.7 68.1 66.3 33.4
GPT-5 82.5 71.0 69.8 70.4 32.1
GPT-5 + RAG 83.6 72.0 75.0 73.5 31.9
@@
ANAH-v2 53.3 11.7 31.0 16.9 46.8
MiniCheck 7B 60.8 25.7 55.4 35.1 43.2
Lynx 8B Instruct v1.1 61.7 37.0 34.8 35.9 42.4
Lynx 70B Instruct 66.1 38.7 54.2 45.1 42.1
@@
LLM-Check (Attn Score) 53.6 59.3 18.3 28.0 47.5
LLM-Check (Hidden Score) 54.7 56.0 19.2 28.6 48.6
ReDeEP 57.1 66.7 20.1 30.8 43.3
InFusE 46.8 38.0 14.4 20.9 52.1
CCP 58.9 48.0 23.8 31.9 44.2
@@
The main evaluation results are shown in Table 5.3. As shown, the most effective methods are zero-shot LLMs, notably GPT-5 + RAG, which achieve balanced accuracy of 83.6%, hallucination F1 of 73.5%, and ranking loss of 31.9%. Although zero-shot methods perform well in coarse-grained unfaithfulness detection, their fine-grained detection performances still face significant challenges.
Figure 4 shows the fine-grained precision and recall of zero-shot LLMs. Across all models, recalls for Ambiguous and Out-Dependent remain consistently low. Nevertheless, compared to GPT-5, GPT-5 + RAG significantly boosts the precision of Out-Dependent while maintaining the recall, showing the effectiveness of using RAG. The low recall for Ambiguous may stem from both the scarcity of Ambiguous examples and the models’ lack of sensitivity to linguistic ambiguity. As for Out-Dependent, we found this class was mainly misclassified as faithful classes (see Figure 7 in Appendix G), indicating that the low recall may be due to the LLM detectors’ unawareness of using external knowledge from their parametric memory.
@@ 6 Conclusion
Acknowledgements
I want to express my sincere gratitude to ZHONG Yang for his careful review and correction of dozens of annotation errors in the dataset, which significantly improved the data quality.
@@ Appendix B Details of Generating the Summaries
Appendix C Special Cases of Annotation
Quotations and Citations. There are many quotations and citations in the collected documents. Inspired by the conventions of academic paper reading, we consider quotations and citations with clear sources to be reliable. Therefore, it is appropriate for the summary to state their contents with confidence. For anonymous citations, we regard them as uncertain facts and consider summaries that do not convey uncertainty as not supported by the documents.
Meta Notes. For some LLMs, there are many meta notes in the generated text, e.g., “*(Note: The mention of Mo Farah’s 2016 gold medal is unrelated to the core information and excluded.)*”. Although these meta notes state some facts, they are not part of the summaries. Therefore, we label these meta notes as No-Fact to exclude them from evaluation.
Appendix D Sequential Dependency of Unfaithfulness
@@
Appendix E Implementation Details of Baselines
@@
Appendix F Zero-shot Detector Details
@@
Appendix G The Confusion Matrix of GPT-5 + RAG